OpenStudy (anonymous):

Choose the Standard Form of the equation y = −x − 6. (4 points) x + y = −6 4x + y = –6 x + 4y = –24 –4x + y = −24

OpenStudy (anonymous):

do you know what standard form of a line is?

OpenStudy (anonymous):

forms of equations of a line: standard: ax+by=c; a, b do not equal 0 slope: -a/b y intercept: c/b point slope: y-y1 = m(x-x1) slope: m y1 and x1 together are the coordinates of a point on the line slope intercept: y=mx+b slope: m y intercept: b

OpenStudy (anonymous):

add x to both sides of the equation: x+y=-6

OpenStudy (anonymous):

Standard form: ax + by = c Your equation: y= -x - 6 All you need to do is carry the x to the left side x+y=-6

OpenStudy (anonymous):

thanks !